# A Self-Supervised Model for Multi-modal Stroke Risk Prediction

### v1

@@ 4.2.2 Imaging heatmaps
Figure 3 shows results from the GradCAM experiment obtained over predicted samples. When inspecting the positive predicted scans (True Positives and False Positives), the model tends to highlight anatomical regions surrounding the lateral ventricles and (periventricular) white matter areas. Such patterns could be associated to white matter hyperintensities, which are known predictors of brain atrophy and age-related brain alterations Grajauskas et al. [2019] and also stroke risk predictors in elderly individuals Park et al. [2019]. In different studies, correlations have been observed between common age-related structural brain changes and brain pathologies Grajauskas et al. [2019], Guo et al. [2017], Shim et al. [2015]. When assessing scans #2 and #4 of the true positive patients in Fig. 3, the activation maps are also showing anatomical regions distant from the lateral ventricles, showing high activations. Supported from literature, those activations could be related to white matter hyperintensities (deep white matter, in this case), often appearing in regions of the brain that are not immediately adjacent to the cortical surface, but commonly located in subcortical white matter or in deep white matter tracts Grajauskas et al. [2019]. Such deep white matter hyperintensities are associated with chronic vascular disease and other chronic pathologies (e.g. multiple sclerosis) Grajauskas et al. [2019]. When evaluating negatively predicted patients (True Negatives and False Negatives), the scans are showing less emphasis on the (periventricular) white matter region but instead highlight areas of the lower brain (cerebellum, posterior brain) and the cortex. We hypothesize that these areas may reflect patterns related to normal aging or normal brain atrophy processes, rather than anomalous brain conditions. Overall, we can hypothesise from these visualizations that the multi-modal SSL stroke risk predictor model focuses on abnormal brain aging patterns for its predictions. We therefore believe that future experiments including brain-age and brain structure-age biomarkers could help enhancing the models predictability, since they have been shown to be associated with overall cardiovascular risk De Lange et al. [2020], clinical outcome in stroke Liew et al. [2023] and overall risk of mortality Cole et al. [2018].
@@
Figure 3: GradCAM-activated brain regions for five patients, categorized as TP (True Positive), TN (True Negative), FP (False Positive), and FN (False Negative). Red (blue) indicates higher (lower) activations.
@@
Limitations. Our study is limited by the use of the UK Biobank, whose demographic characteristics may not fully represent the diversity of global populations, potentially impacting the model’s generalizability and clinical utility. Future research should validate our approach using more diverse external datasets to improve applicability. Additionally, our test set was constrained by the limited availability of pre-stroke imaging samples, as most stroke datasets focus on post-onset cases. Finally, the heterogeneity in the time between imaging and stroke onset in the UK Biobank could influence model performance, necessitating further experiments to disentangle these effects. Future work could also include improving model efficiency by testing further architectures and techniques to reduce model parameters (e.g. network pruning).
@@ 5 Conclusion
We hereby present an SSL model integrating diverse data modalities for predicting stroke risk. The model’s performance is compared against state-of-the-art (self-)supervised models employing both unimodal and multi-modal data, including tabular and imaging datasets. A comprehensive set of experimental settings is utilized, encompassing different subgroupings of tabular features—such as clinical data, brain IDPs, and lesion IDPs—as well as various training regimes that combine pre-training and fine-tuning based on data modality.
@@
Our results demonstrate that the CLIP model on multi-modal data, combined with an ITM loss, outperforms single-modality alternatives.The CLIP+ITM model surpasses the self-supervised tabular (image) data SCARF (SimCLR) model by 2.6% (2.6%) in ROC-AUC, and by 3.3% (5.6%) in balanced accuracy terms. Our framework also demonstrated an AUROC improvement of 0.93% and 7.6% balanced accuracy from the best multi-modal supervised method. Additionally, the proposed model produces well-aligned multi-modal representations in a common, data modality-independent space, which is unattainable with unimodal tabular or imaging data models. Thus, CLIP-ITM effectively leverages complementary and synergistic information from diverse data modalities.
@@
Using interpretable GradCAM heatmaps, we identified activated brain regions commonly associated with brain aging, stroke risk, and clinical outcomes. On one hand, the activated areas indicate that the model primarily focuses on deep and periventricular white matter hyperintensities for predicting positive samples, which may be more common and extensive in patients identified as at risk for stroke. On the other hand, the prediction of negative samples highlights the cerebellum, posterior brain regions and cortical areas. These results demonstrate the model’s capacity to extract task-specific features linked to stroke risk, which are well-supported by existing literature.
@@
In conclusion, we propose a robust self-supervised multi-modal learning approach for stroke risk prediction. Our model offers a strong foundation for future studies that aim to integrate multiple data modalities into prediction models.
Acknowledgement
@@
This research has been conducted using the UK Biobank resource under application number 81959 and grants NHLBIR21HL156174, NHLBIK24HL136852, and K24HL136852. This project was supported by the grant #2023-N-306 of the 1st Joint Call of the Swiss Data Science Center (SDSC) and the Strategic Focus Area “Personalized Health and Related Technologies (PHRT)” of the ETH Domain (Swiss Federal Institutes of Technology). N.D. is partially supported by the ETH AI Center postdoctoral fellowship. B.M., N.D., and E.D. acknowledge support of the Helmut-Horten-Foundation.

### v2

@@ 4.2.2 Imaging heatmaps and tabular feature importance
Figure 5(b) shows the results from the GradCAM experiment obtained over predicted samples. When inspecting the positive predicted scans (True Positives and False Positives), the model tends to highlight anatomical regions surrounding the lateral ventricles and (periventricular) white matter areas. Such patterns could be associated to WMHs, which are known predictors of brain atrophy and age-related brain alterations [Grajauskas2019] and also stroke risk predictors in elderly individuals [park2019white]. In different studies, correlations have been observed between common age-related structural brain changes and brain pathologies [Grajauskas2019, Guo2017, Shim2015]. In the True Positive patients class, the views of scans #2 and #4 in Fig. 5(b) with overlaid activation maps are highlighting distant anatomical regions from the lateral ventricles. Supported from literature, those activations could be related to deep WMHs, often appearing in brain regions that are not immediately adjacent to the cortical surface, but commonly located in subcortical white matter or in deep white matter tracts [Grajauskas2019]. Such DWMHs are associated with chronic vascular disease and other chronic pathologies (e.g. multiple sclerosis) [Grajauskas2019]. When looking at negatively predicted patients (True Negatives and False Negatives), the scans are putting less emphasis on the periventricular white matter region but instead highlight areas of the lower brain (cerebellum, posterior brain) and the cortex. We hypothesize that these areas may reflect patterns related to normal aging or normal brain atrophy processes, rather than anomalous brain conditions. Overall, we can hypothesize from these visualizations that the multi-modal SSL stroke risk predictor model focuses on pathological lesions for its predictions. We therefore believe that further experiments including brain age and brain structural age biomarkers could help enhancing the model’s predictability, since they have been shown to be associated with overall cardiovascular risk [de2020multimodal], clinical outcome in stroke [liew2023association] and overall risk of mortality [cole2018brain].
(a) Tabular feature importance.
(b) GradCAM-activated brain regions
@@
Figure 5: Feature importance analysis for model interpretability with anatomical T2-FLAIR scans and clinical records. (A) Tabular feature importance for image-derived (orange) and clinical features (blue) from integrated gradients analysis. (B) GradCAM-activated brain regions for five patients, categorized as TP (True Positive), TN (True Negative), FP (False Positive), and FN (False Negative). Red (blue) indicates higher (lower) activations.
Figure 6: Boxplot of normalized GradCAM activations of white matter hyperintensities masks over whole brain. The distributions are constructed using per-scan average activation values. $p$-value is shown in bold.
Figure 6 shows the distributions of normalized average GradCAM activations across WMHs masks over the whole brains. As evidenced by the figure, stroke patients are showing a significantly higher mean activation in WMH regions than for overall brains compared to the healthy group, suggesting that GradCAM has identified more pronounced activations in regions correlating with stroke development [Grajauskas2019]. A Mann-Whitney U test was performed to compare the normalized mean GradCAM activations between stroke and healthy groups and shows a statistically significant difference (U = 1403, $p$-value: 1.35 x 10-2), indicating distinct normalized mean activation levels between the two groups. As these activations highlight areas of interest, particularly in WMH lesions, it is likely that the model is focusing on relevant pathological regions when identifying stroke-related features. This significant difference underscores the model’s sensitivity to stroke markers, especially when WMH lesions are isolated from the whole brain MRI, indicating reliable discrimination between stroke and healthy brain images based on observed activations.
@@
Limitations. Our study is limited by the use of the UK Biobank, whose demographic characteristics may not fully represent the diversity of global populations, potentially impacting the model’s generalizability and clinical utility. Future research should validate our approach using more diverse external datasets to improve applicability. Additionally, our test set was constrained by the limited availability of pre-stroke imaging samples, as most stroke datasets focus on post-onset cases. Therefore, it is not trivial to define a standardized time window where individuals would be monitored during the pre-stroke onset period. However, from the scan acquisition to either the stroke onset or the second healthy scan acquisition, our algorithm spans between 26 and 29 months on average for all participants. Moreover, there is some variability in the model performance with respect to the time between imaging and stroke onset, even though there is no clear trend showing that it is negatively affecting the model performance. Future work could also include improving model efficiency by testing further architectures and techniques such as network pruning [Wang2024] to reduce model parameters. Finally, analyzing saliency maps (e.g., Grad-CAMs) using calibrated models or advanced techniques (such as Score-CAM or Opti-CAM [Wang2020ScoreCAM, Zhang2024OptiCAM, Moller2020ReliableSaliency]) could yield more precise insights into imaging biomarkers and improve voxel-level localization of brain regions that most strongly influence the model’s decisions, ultimately enhancing our understanding of stroke risk predictors.
@@ 5 Conclusion
We hereby present a multi-modal foundation model integrating diverse data modalities for predicting stroke risk. The model performance is compared against state-of-the-art (self-) supervised models employing both uni-modal and multi-modal data, including tabular and imaging datasets. A comprehensive set of experimental settings is utilized, encompassing different subgroupings of tabular features - such as clinical records, brain IDPs, and lesion IDPs - as well as various training regimes that combine pre-training and fine-tuning based on data modality.
@@
Our results demonstrate that the CLIP model on multi-modal data, combined with an ITM loss, outperforms single-modality alternatives. Our method surpasses the self-supervised tabular (image) data SCARF (SimCLR) model by 2.1% (2.8%) in ROC-AUC, and by 10.6% (12.8%) in balanced accuracy terms. It also demonstrated a 0.93% AUROC and 7.6% accuracy improvements when compared with the best multi-modal supervised method. Additionally, the proposed model produces well-aligned multi-modal representations in a common, data modality-independent space, which is unattainable with uni-modal tabular or imaging data models. Thus, our method effectively leverages complementary and synergistic information from diverse data modalities.
@@
Using interpretable heatmaps, we identified activated brain regions commonly associated with brain aging, stroke risk, and clinical outcomes. On one hand, the activated areas indicate that the model primarily focuses on deep and periventricular WMHs for predicting positive samples, which may be more common and extensive in patients identified as at risk for stroke. On the other hand, the prediction of negative samples highlights the cerebellum, posterior brain regions and cortical areas. These results demonstrate the model capacity to extract task-specific features linked to stroke risk, which are well-supported by existing literature.
@@
In conclusion, we propose a robust multi-modal foundation model for stroke risk prediction. Our model lays the groundwork for future studies aiming to integrate multiple data modalities into prediction models.
6 Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
7 Non-standard abbreviations
  • PWMHs, DWMHs, WMHs: Periventricular, Deep, white matter hyperintensities
  • SSL: Self-supervised learning
  • MRI: Magnetic Resonance Imaging
  • CLIP: Contrastive Language-Image Pre-training
  • ITM: Image-Tabular Matching
  • CNN: Convolutional Neural Networks
  • UKB: UK Biobank
  • ICD: International Classification of Disease
  • IDPs: Image-Derived Phenotypes
  • MLP: Multi-Layer Perceptron
  • UMAP: Uniform Manifold Approximation and Projection
  • GradCAM: Gradient-weighted Class Activation Mapping
  • ROC: Receiver Operating Characteristic
  • AUC: Area Under the Curve
8 Funding Statement and Acknowledgments
@@
This research has been conducted using the UK Biobank resource under application number 81959 and grants NHLBI R21HL156174, K24HL136852, R21HL167173. The project was supported by the grant #2023-N-306 of the 1st Joint Call of the Swiss Data Science Center (SDSC) and the Strategic Focus Area “Personalized Health and Related Technologies (PHRT)” of the ETH Domain (Swiss Federal Institutes of Technology) and DataSpectrum4CVD and the grant from PHRT Project #2022-812 / SDSC C22-15P. BM, ND, and EDLR acknowledge support of the Helmut-Horten-Foundation.
9 Data Availability Statement
The UK Biobank dataset is semi-private. You can apply for an academic access here: access UKBB. The code repository for this project is available at the following link: SSMSRPM.
[…]